Nature Place operates a commercial plant nursery, where it propagates plants for garden centers throughout the region. Nature Place has $5,100,000 in assets. Its yearly fixed costs are $600,000 and the variable costs for the potting soil, container, label, seedling and labor for each gallon-size plant total $1.25. Nature Place's volume is currently 500,000 units. Competitors offer the same plants, at the same quality to garden centers for $3.50 each. Garden centers then mark them up to sell to the public for $9 to $12 depending on the type of plant.

Requirements

  1. Nature Paces owners want to earn a 11% return on the company' assets. What is Nature Place's target full product cost?
  2. Given Nature Place's current costs. Will its owners be able to achieve their target profit?
  3. Assume Nature Place has identified ways to cut its variable costs to $1.10 per unit what is its new target fixed cost? Will this decrease in variable costs allow the company to achieve its target profit?
  4. Nature Place started an aggressive advertising campaign strategy to differentiate its plants from those grown by other nurseries. Monrovia Plants made this strategy work so Nature Pace has decided to try it, too. Nature Place does not expect volume to be affected, but it hopes to gain more control over pricing. If Nature Place has to spend $80,000 this year to advertise, and its variable costs continue to be $1.10 per unit, what will its cost-plus price be? Do you think Nature Place will be able to sell its plants to garden centers at the cost-plus price? Why or why not?
Answer #1

Ans 1) Nature place Expected Return

  • Expected Return on Assets= Assets total value* rate= $ 51,00,000*11/100=5,61,000,
  • Per Unit Expected Return= $ 1.122 (Expected Profit per unit)
  • Total cost of Nature Place

-Fixed Cost                                       = $ 6,00,000

-Variable Cost(1.25*5,00,000)    = $ 6,25,000

                               Total Cost    = $ 12,25,000

       Per Unit Cost is $ 12,25,000/500000=$2.45

Per  Unit Selling Price= $2.45+$ 1.122 =$ 3.572

*Per Unit Selling Price should be $ 3.572 including the 11% profit margin.

Ans 2) No, as the current selling price per unit is less than the price at which it should be sold i.e.

Expected Price= $ 3.572

Current Selling Price= $ 3.50.

the difference of per unit price is $ 0.072 making a difference of $ 36,000 shortfall of the target.

Ans 3) If nature cost Reduces it variable cost to $1.10 then the new fixed cost can be

          Variable cost= 5,00,000*$1.10=5,50,000

         New fixed Cost= $12,25,000(Total Cost) - $5,50,000(Variable Cost)-Shortfall Profit

        New Fixed cost = $ 6,39,000.

So the owner is having the option to increase the cost from $6,00,000 to $ 6,39,000   with a return of 11%.Yes the owner will be able to achive its 11% of profit.

Ans 4 ) Nature Fresh is having a excess liquidity of $ 39,000 with them which they have to adjust agaist the advertisement expense of $ 80,000 as it creates a shortfall of $ 41,000 because of which the nature place will not be able to sell its plants to garden place at cost-plus price. Its calculation will be as follows :

  • Expected Return on Assets= Assets total value* rate= $ 51,00,000*11/100=5,61,000,
  • Per Unit Expected Return= $ 1.122 (Expected Profit per unit)
  • Total cost of Nature Place

-Fixed Cost                                       = $ 6,00,000

-Variable Cost(1.10*5,00,000)    = $ 5,50,000

-Advertisment Cost                        = $ 80,000

                               Total Cost    = $ 12,30,000

       Per Unit Cost is $ 12,30,000/500000=$2.46

      Selling Price Per Unit = cost + profit

                              = $2.46 + $ 1.122 = $ 3.582

                              =Selling price + Expected Price

                               = $ 3.582- $ 1.122= $ 0.082

         So there is a difference of $0.082*500000= $ 41,000.